solve the following system of equation 2x-3y=13 and 5x+2y=4
Answer by stanbon(75887)

solve the following system of equation
2x-3y=13
5x+2y=4
------
Multiply 1st by 5
Multiply 2nd by 2
-------
10x - 15y = 65
10x + 4y = 8
---
Subtract and solve for "y":
19y = -57
y = -3
----
Solve for "x":
2x - 3y = 13
2x + 9 = 13
2x = 4
x = 2
-------
Cheers,
Stan H.